Session: Nov 1999 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Bondi at Tue Nov 16 23:17:39 MET 1999) Minor failures

Effelsberg (A. Kraus at Wed Nov 17 08:40:31 MET 1999) Success

Torun (K.Borkowski at Wed Nov 17 09:29:15 MET 1999) Success
No known problems.

Onsala (A. Polatidis at Wed Nov 17 12:50:50 MET 1999) Minor failures
We had to stow the telescope from 07:53 UT due to high winds.

Westerbork (Tony Foley at Wed Nov 17 13:28:27 MET 1999) Minor failures
Data lost from 19:00 to 21:07 (day 319) as the frontend frequency control software was out. There was also a huge phase jump in several frontends from about 10:30 to 19:00 (day 320), causing a large drop in array sensitivity. The problem was a 50ohm terminator for an LO signal. 12 telescopes in array.

Noto (P.Leto at Wed Nov 17 17:50:19 MET 1999) Success

Jodrell (Alastair Gunn at Fri Nov 19 12:27:04 MET 1999) Minor failures
Few minutes late on first source. No other known problems.

Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:01:48 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: ey004d.antab. Read the comments in the file to get further information. (Rem : For Ef, bad Tsys valules at the end of the exp).